# On absolutely continuous spectrum for one-channel unitary operators

### v1

@@ 5. Perturbation of periodic scattering zipper
Let ${\mathcal{U}}$ be a one-channel scattering zipper fulfilling assumptions (A1), (A2), meaning that $b_{n}\neq 0,{\mathfrak{b}}_{n}\neq 0$ for asll $n$. Then, the vector $e_{(0,-)}$ is cyclic, meaning
@@
Note that for scattering zippers, the vectors $e_{(n,\pm)}$ form an ortonormal basis of $\ell^{(}{\mathbb{G}})$. By induction we will prove that for all $n\in{\mathbb{Z}}_{+}$, $e_{(n,-)},e_{(n,+)}$ are in the ${\mathcal{U}}$-cyclic space ${\mathcal{Z}}$ generated by $e_{(0,-)}$.
@@
${\mathcal{U}}e_{(n,-)}\,=\,{\mathcal{W}}\left(a_{n}e_{(n,-)}+c_{n}e_{(n,+)}\right)\,=\,\underbrace{a_{n}{\mathcal{W}}e_{(n,-)}\,+\,c_{n}{\mathfrak{a}}_{n+1}e_{(n,+)}}_{\in{\mathcal{Z}}}\,+\,c_{n}{\mathfrak{b}}_{n+1}e_{(n+1,-)}$
Note, as $b_{n}\neq 0$ also implies $c_{n}\neq 0$, and we always have ${\mathfrak{b}}_{n+1}\neq 0$. Hence, we see that $e_{(n+1,-)}\in{\mathcal{Z}}$. Now, with the induction hypothesis, ${\mathcal{W}}e_{(n+1,-)}\,=\,{\mathfrak{b}}_{n+1}e_{(n,+)}+{\mathfrak{d}}_{n+1}e_{(n+1,-)}\in{\mathcal{Z}}$, and therefore
@@
Furthermore, note that for ${\mathcal{U}}$ and ${\mathcal{U}}_{\omega}$ we have that the set ${\mathcal{A}}$ as defined in (1.13), (1.14) is empty. Hence, the measure $\nu$ (or $\nu^{(u)}$) as in Theorem 2 is equal to zero. As in Section 1.4 we have the transfer matrix over a period $T_{z}=T_{z,[1,p]}$ and recall $\Sigma=\{z\in{\mathbb{U}}(1)\,:\,|\operatorname{{\rm Tr}}T_{z}|<2\}$. We now show that $\Sigma$ is always a subset of the absolutely continuous spectrum (almost surely).
@@
There is a set $\Omega^{\prime}\subset\Omega$ or probability one, such that for all $\omega\in\Omega^{\prime}$, the spectral measure $\mu_{\omega}$ of ${\mathcal{U}}_{\omega}$ at the vector $e_{(0,-)}$ is purely absolutely continuous in $\Sigma$ and $\Sigma\subset\operatorname{{\rm supp}}\mu_{\omega}$. We also find that $\mu$ is purely absolutely continuous in $\Sigma$ and $\Sigma\subset\operatorname{{\rm supp}}\mu$.
@@
By Proposition A.4, for $z\in\Sigma$, there is $\varphi_{z},\theta_{z}\in{\mathbb{R}}$, $M_{z}\in{\rm GL}(2,{\mathbb{C}})$ such that
@@
which is a perturbation of $T_{z}^{n}T_{z,0}$. Now, as $V_{n},\widehat{V}_{n},W_{n},\widehat{W}_{n}$ are unitary matrices, their norms are all igual to one. Thus, with assumption (C3), the matrices $\|T_{z,n}\|$, $\|\widehat{T}_{z,n}\|$ are uniformly bounded by some constant $C$. Possibly increasing the constant $C$ above, we may use the same bound.
@@
uniformly in $z\in{\mathbb{U}}(1)$. Now, as we have
$\widetilde{T}_{z,n}-T_{z}$ $=\widetilde{T}_{z,n}-T_{z}\,=\,\prod_{k=1}^{p}\widehat{T}_{z,(n-1)p+k}\,-\,\prod_{k=1}^{p}T_{z,(n-1)p+k}$
$=\sum_{k=1}^{p}\left(\prod_{l=k+1}^{p}\widehat{T}_{z,(n-1)p+k}\,\big{[}\widehat{T}_{z,(n-1)p+k}-T_{z,(n-1)p+k}\big{]}\prod_{l=1}^{k-1}T_{z,(n-1)p+l}\right)$
@@
$\|\widetilde{T}_{z,n}-T_{z}\|\,\leq\,C^{p-1}\sum_{k=1}^{p}\|\widehat{T}_{z,(n-1)p+k}-T_{z,(n-1)p+k}\|.$
Using independence and the expected matrices of the $\hat{T}_{z,n}$ one gets similarly
$\big{\|}{\mathbb{E}}\big{(}\widetilde{T}_{z,n}\big{)}-T_{z}\big{\|}\,\leq\,C^{p-1}\sum_{k=1}^{p}\big{\|}{\mathbb{E}}\big{(}\widehat{T}_{z,(n-1)p+k}\big{)}-T_{z,(n-1)p+k}\big{\|}\,\quad\text{for $n\geq 1$}\,.$
@@
Then, using (5.1), (5.2) and the fact that $\widetilde{T}_{z,0}$ is uniformly bounded, we obtain
$\sum_{n=0}^{\infty}\left(\big{\|}{\mathbb{E}}(S_{z,n})\big{\|}\,+\,{\mathbb{E}}\big{(}\|S_{z,n}\|^{2}\big{)}\right),\,<\,\tilde{C}<\infty$ (5.3)
@@
$\quad+\,{\mathbb{E}}\Big{(}2\|{\vec{v}}_{n}\|^{2}\,\|S_{z,n}{\vec{v}}_{n}\|^{2})+4\,\Re e\big{(}{\,{\vec{v}}_{n}}^{*}R_{z}^{*}{\mathbb{E}}(S_{z,n})\big{)}{\vec{v}}_{n}\,\|{\vec{v}}_{n}\|^{2}\,\Big{)}$
@@
$\leq\,{\mathbb{E}}(\|{\vec{v}}_{n}\|^{4})\exp\Big{(}c\;\big{(}\big{(}\|{\mathbb{E}}(S_{z,n})\|\,+\,{\mathbb{E}}(\|S_{z,n}\|^{2})\Big{)}$
@@
We note that by Proposition A.4 we find for $z\in{\mathbb{U}}(1)$ that $|\operatorname{{\rm Tr}}T_{z}|=2\Leftrightarrow\tfrac{(\operatorname{{\rm Tr}}T_{z})^{2}}{\det T_{z}}=4$. Moreover, $z\mapsto\tfrac{(\operatorname{{\rm Tr}}T_{z})^{2}}{\det T_{z}}$ is analytic for $z\neq 0$. Therefore, if this would be equal to 4 on an infinite set within the unit circle ${\mathbb{U}}(1)$, then the function would have to be identically $4$, and so would be $(\operatorname{{\rm Tr}}z^{p}T_{z})/\det(z^{p}T_{z})$. Now, using (1.19), (1.20) one finds for $n\geq 1$
@@
Let us first show Claim 1: For the unique formal solutions (uniqueness because needs to have the correct boundary condition) of ${\mathcal{U}}\Psi=z\Psi$ and ${\mathcal{W}}\Phi=\Psi$ we have that
$\begin{pmatrix}\Phi_{(n,+)}\\ \Psi_{(n,+)}\end{pmatrix}\,=\,T_{[0,n]}\begin{pmatrix}1\\ 1\end{pmatrix}\,,\quad\begin{pmatrix}\Phi_{(n,+)}\\ \Psi_{(n,+)}\end{pmatrix}\,=\,T_{z,n}^{\sharp}\,\begin{pmatrix}\Psi_{(n,-)}\\ \Phi_{(n,-})\end{pmatrix}.$
We have the following equivalences
@@
$\begin{pmatrix}\Phi_{(np,+)}\\ \Psi_{(np,+)}\end{pmatrix}\,=\,T_{z}^{p}{[0,n]}\begin{pmatrix}1\\ 1\end{pmatrix}\,=\,T_{z}^{n}T_{z,0}\begin{pmatrix}1\\ 1\end{pmatrix},$
@@
And finally we remark that there is no spectrum where $|\operatorname{{\rm Tr}}T_{z}|<2$.
@@
and as $z$ is not n eigenvalue of ${\mathcal{U}}$, Lemma 5.4 implies $y_{z}\neq 0$.Thus, $|y_{z}|>\varepsilon>0$ for some $\varepsilon>0$ and all $z\in e^{i[a,b]}$. Thus, for $f\in C({\mathbb{U}}(1))$ supported in $e^{i[a,b]}$ we have using Theorem 1
@@
Note that by Lemma 5.1 the spectrum of ${\mathcal{U}}$ has multiplicity 1 everywhere and is given by the support and measure types of $\mu$. Then Lemmata 5.3, 5.4 and 5.5 imply that there is no essential spectrum in the interior222interior with respect to the topology on the unit circle ${\mathbb{U}}(1)$. of $\{z\in{\mathbb{U}}(1)\,:\,|\operatorname{{\rm Tr}}T_{z}|\geq 2\}$. So Lemma 5.2 now shows
@@
Together with Lemma 5.4 we deduce that the spectrum is purely absolutely continuous within $\overline{\Sigma}$. Apart from $\overline{\Sigma}$ the spectrum contains only a finite number of eigenvalues outside $\overline{\Sigma}$. Finally, as the Hilbert space is infinite dimensional and the spectrum has multiplicity one, a finite set of eigenvalues can not be the whole spectrum. Thus, $\Sigma\neq\emptyset$. This finishes the proof of Proposition 1.5.
Then, the fact that ${\mathcal{U}}_{\omega}$ is almost surely a compact perturbation of ${\mathcal{U}}$ by condition (C2) shows that as a set, the essential spectrum of ${\mathcal{U}}_{\omega}$ is also given by $\overline{\Sigma}$, and Lemma 5.2 now shows Theorem 3.
@@ Appendix A Facts of linear algebra
$U=\begin{pmatrix}A&B\\ C&D\end{pmatrix}\,\in\,{\mathbb{U}}(n)$
be a unitary $n\times n$ matrix split in blocks of sizes $n_{1},n_{2}$, meaning $A\in{\mathbb{C}}^{n_{1}\times n_{1}}$, $B\in{\mathbb{C}}^{n_{1}\times n_{2}}$, $C\in{\mathbb{C}}^{n_{2}\times n_{1}}$, $D\in{\mathbb{C}}^{n_{2}\times n_{2}}$. Let $V\in{\mathbb{U}}(n_{2})$. Then, the following holds.
@@
    Moreover, if $B$ has trivial kernel then $D-V$ is invertible (for any $V\in{\mathbb{U}}(n_{2})$).
@@
    Assume that $V\in{\mathbb{U}}(n_{2})$ and that $D-V$ is invertible. Then, the Schur complement
    $A-B(D-V)^{-1}C\,\in\,{\mathbb{U}}(n_{1})$
@@
    $\begin{pmatrix}\mathbf{I}&\mathbf{0}\end{pmatrix}\begin{pmatrix}A&B\\ C&D-V\end{pmatrix}^{-1}\begin{pmatrix}\mathbf{I}\\ \mathbf{0}\end{pmatrix}\,\in\,{\mathbb{U}}(n_{1})$
@@
    More general, let $V,\,W\in{\mathbb{C}}^{n_{2}\times n_{2}}$ such that $W^{*}V=\mathbf{I}_{n_{2}}$, and $C-V$ and $D-W$ are invertible, then
    $(A-B(D-W)^{-1}C)^{*}(A-B(D-V)^{-1}C)\,=\,\mathbf{I}_{n_{1}}.$
@@
Let $|z|=1$. First note that if $(D-V)\varphi=\mathbf{0}$, then
$\|\varphi\|^{2}=\left\|U\begin{pmatrix}\mathbf{0}\\ \varphi\end{pmatrix}\right\|^{2}=\|B\varphi\|^{2}+\|D\varphi\|^{2}=\|B\varphi_{2}\|+\|V\varphi\|^{2}=\|B\varphi\|^{2}+\|\varphi\|^{2},$
@@
On the other hand, let $\begin{pmatrix}A&B\\ C&D-V\end{pmatrix}\begin{pmatrix}\varphi_{1}\\ \varphi_{2}\end{pmatrix}=\mathbf{0}$, then
$\|\varphi_{1}\|^{2}+\|\varphi_{2}\|^{2}=\left\|\begin{pmatrix}\varphi_{1}\\ \varphi_{2}\end{pmatrix}\right\|^{2}=\left\|U\begin{pmatrix}\varphi_{1}\\ \varphi_{2}\end{pmatrix}\right\|^{2}=\left\|\begin{pmatrix}\mathbf{0}\\ V\varphi_{2}\end{pmatrix}\right\|^{2}=\|\varphi_{2}\|^{2}\;.$
Therefore, $\varphi_{1}=\mathbf{0}$ and $D\varphi_{2}=V\varphi_{2}$. Hence,
$\ker\begin{pmatrix}A&B\\ C&D-V\end{pmatrix}\,=\,\left\{\begin{pmatrix}\mathbf{0}\\ \varphi\end{pmatrix}\,:\,\varphi\in\ker(D-V)\;\right\}\;.$
For part b) note that the first part of b) follows from c) Using the Schur complement identity
$\begin{pmatrix}\mathbf{I}&\mathbf{0}\end{pmatrix}\begin{pmatrix}A&B\\ C&D-V\end{pmatrix}^{-1}\begin{pmatrix}\mathbf{I}\\ \mathbf{0}\end{pmatrix}\,=\,(A-B(D-V)^{-1}C)^{-1}\;,$
@@
and using $W^{*}V=\mathbf{I}$ we obtain
$(A-B(D-W)^{-1}C)^{*}(A-B(D-V)^{-1}C)$
$=A^{*}A-A^{*}B(D-V)^{-1}C-C^{*}(D^{*}-W^{*})^{-1}B^{*}A+C^{*}(D^{*}-W^{*})^{-1}B^{*}B(D-V)^{-1}C$
$=A^{*}A+C^{*}D(D-V)^{-1}C+C^{*}(D^{*}-W^{*})^{-1}D^{*}C+C^{*}(D^{*}-W^{*})^{-1}(\mathbf{I}-D^{*}D)(D-V)^{-1}C$
$=A^{*}A+C^{*}(D^{*}-W^{*})^{-1}\big{(}(D^{*}-W^{*})D+D^{*}(D-V)+\mathbf{I}-D^{*}D\big{)}(D-V)^{-1}C$
$=A^{*}A+C^{*}(D^{*}-W^{*})^{-1}\big{(}(D^{*}-W^{*})(D-V)\big{)}(D-V)^{-1}C=A^{*}A+C^{*}C=\mathbf{I}\;.$
@@
For $T\in{\mathbb{U}}(1,1)$ we have $|\det T|=1$ and $\tfrac{(\operatorname{{\rm Tr}}T)^{2}}{\det T}\geq 0$. In particular, there is $\chi\in{\mathbb{R}}$ and $\lambda\in{\mathbb{C}}$, such that the eigenvalues of $T$ are given by $e^{i\chi}\lambda$ and $e^{i\chi}\lambda^{-1}$ with $\lambda+\lambda^{-1}\in{\mathbb{R}}$.
@@
then $T^{*}GT=G$ which means $-1=\det(G)=\det(G)|\det T|^{2}$ implying $|\det(T)|=1$. Moreover, we get $|a|^{2}-|c|^{2}=a=|b|^{2}-|d|^{2}$ and $\bar{a}b=\bar{c}d$. On the other hand, $T^{*}GTG=G^{2}=I$ and therefore $TGT^{*}G=I$ giving $TGT^{*}=G$, which gives $|a|^{2}-|b|^{2}=1=|d|^{2}-|c|^{2}$. Hence, $|a|=|d|=\cosh(\gamma)$ and $|b|=|c|=\sinh(\gamma)$ for some $\gamma\geq 0$.
@@
$a=\cosh(\gamma)e^{i\varphi_{1}},\;b=\sinh(\gamma)e^{i\varphi_{2}},\;c=\sinh(\gamma)e^{i\varphi_{3}},\;d=\cosh(\gamma)e^{i\varphi_{4}}$
@@
$e^{-i\varphi_{1}+i\varphi_{2}}\,=\,e^{-i\varphi_{3}+i\varphi_{4}}\quad\text{implying}\quad e^{i\varphi_{1}+i\varphi_{4}}=e^{i\varphi_{2}+i\varphi_{3}}$

### v2

@@ 5. Perturbation of periodic scattering zipper
Let ${\mathcal{U}}$ be a one-channel scattering zipper fulfilling assumptions (A1), (A2), meaning that $b_{n}\neq 0,{\mathfrak{b}}_{n}\neq 0$ for all $n$. Then, the vector $e_{(0,-)}$ is cyclic, meaning
@@
Note that for scattering zippers, the vectors $e_{(n,\pm)}$ form an orthonormal basis of $\ell^{2}({\mathbb{G}})$. By induction we will prove that for all $n\in{\mathbb{Z}}_{+}$, $e_{(n,-)},e_{(n,+)}$ are in the ${\mathcal{U}}$-cyclic space ${\mathcal{Z}}$ generated by $e_{(0,-)}$.
@@
${\mathcal{U}}e_{(n,-)}\,=\,{\mathcal{W}}\left(a_{n}e_{(n,-)}+c_{n}e_{(n,+)}\right)\,=\,\underbrace{a_{n}{\mathcal{W}}e_{(n,-)}\,+\,c_{n}{\mathfrak{a}}_{n+1}e_{(n,+)}}_{\in{\mathcal{Z}}}\,+\,c_{n}{\mathfrak{b}}_{n+1}e_{(n+1,-)}.$
Note, $c_{n}\neq 0$ (as $b_{n}\neq 0$) and ${\mathfrak{b}}_{n+1}\neq 0$ by assumption. Hence, we see that $e_{(n+1,-)}\in{\mathcal{Z}}$. Now, with the induction hypothesis, ${\mathcal{W}}e_{(n+1,-)}\,=\,{\mathfrak{b}}_{n+1}e_{(n,+)}+{\mathfrak{d}}_{n+1}e_{(n+1,-)}\in{\mathcal{Z}}$, and therefore
@@
Furthermore, note that for ${\mathcal{U}}$ and ${\mathcal{U}}_{\omega}$ we have that the set ${\mathcal{A}}$ as defined in (1.13), (1.14) is empty. Hence, the measure $\nu$ (or $\nu^{(u)}$) as in Theorem 2 is equal to zero. As in Section 1.4 we have the transfer matrix over a period $T_{z}=T_{z,[1,p]}$ and recall $\Sigma=\{z\in{\mathbb{U}}(1)\,:\,|\operatorname{{\rm Tr}}T_{z}|<2\}$. We now show that $\Sigma$ is always a subset of the absolutely continuous spectrum (almost surely). This is the main part of the proof of Theorem 3 and Proposition 1.8.
@@
There is a set $\Omega^{\prime}\subset\Omega$ of probability one, such that for all $\omega\in\Omega^{\prime}$, the spectral measure $\mu_{\omega}$ of ${\mathcal{U}}_{\omega}$ at the vector $e_{(0,-)}$ is purely absolutely continuous in $\Sigma$ and $\Sigma\subset\operatorname{{\rm supp}}\mu_{\omega}$. We also find that $\mu$ is purely absolutely continuous in $\Sigma$ and $\Sigma\subset\operatorname{{\rm supp}}\mu$.
@@
By Proposition A.4, for $z\in\Sigma$, there is $\varphi_{z},\theta_{z}\in{\mathbb{R}}$, (not necesarily equal) and $M_{z}\in{\rm GL}(2,{\mathbb{C}})$ such that
@@
which is a perturbation of $T_{z}^{n}T_{z,0}$. Now, as $V_{n},\widehat{V}_{n},W_{n},\widehat{W}_{n}$ are unitary matrices, their norms are all equal to one. Thus, with assumption (C3), the matrices $\|T_{z,n}\|$, $\|\widehat{T}_{z,n}\|$ are uniformly bounded by some constant $C$. Possibly increasing the constant $C$ above, we may use the same bound.
@@
uniformly in $z\in{\mathbb{U}}(1)$. We have
$\widetilde{T}_{z,n}-T_{z}\,=\,\prod_{k=1}^{p}\widehat{T}_{z,(n-1)p+k}\,-\,\prod_{k=1}^{p}T_{z,(n-1)p+k}$
$=\sum_{k=1}^{p}\left(\left(\prod_{l=k+1}^{p}\widehat{T}_{z,(n-1)p+l}\right)\,\left(\widehat{T}_{z,(n-1)p+k}-T_{z,(n-1)p+k}\right)\left(\prod_{l=1}^{k-1}T_{z,(n-1)p+l}\right)\right),$ (5.3)
@@
$\|\widetilde{T}_{z,n}-T_{z}\|\,\leq\,C^{p-1}\sum_{k=1}^{p}\|\widehat{T}_{z,(n-1)p+k}-T_{z,(n-1)p+k}\|.$ (5.4)
Replacing $\widetilde{T}_{z,n}$ and $\widehat{T}_{z,n}$ with their expectations in (5.3) and using independence leads to
$\big{\|}{\mathbb{E}}\big{(}\widetilde{T}_{z,n}\big{)}-T_{z}\big{\|}\,\leq\,C^{p-1}\sum_{k=1}^{p}\big{\|}{\mathbb{E}}\big{(}\widehat{T}_{z,(n-1)p+k}\big{)}-T_{z,(n-1)p+k}\big{\|}\,\quad\text{for $n\geq 1$}\,.$ (5.5)
@@
Then, using (5.1), (5.2), (5.4), (5.5) and the fact that $\widetilde{T}_{z,0}$ is uniformly bounded, we obtain
$\sum_{n=0}^{\infty}\left(\big{\|}{\mathbb{E}}(S_{z,n})\big{\|}\,+\,{\mathbb{E}}\big{(}\|S_{z,n}\|^{2}\big{)}\right)\,<\,\tilde{C}<\infty,$ (5.6)
@@
$\quad+\,{\mathbb{E}}\Big{(}2\|{\vec{v}}_{n}\|^{2}\,\|S_{z,n}{\vec{v}}_{n}\|^{2})+4\,\Re e\big{(}{\,{\vec{v}}_{n}}^{*}R_{z}^{*}{\mathbb{E}}(S_{z,n}){\vec{v}}_{n}\big{)}\,\|{\vec{v}}_{n}\|^{2}\,\Big{)}$
@@
$\leq\,{\mathbb{E}}(\|{\vec{v}}_{n}\|^{4})\exp\Big{(}c\;\big{(}\big{(}\|{\mathbb{E}}(S_{z,n})\|\,+\,{\mathbb{E}}(\|S_{z,n}\|^{2})\Big{)},$
@@
We note that by Proposition A.4 we find for $z\in{\mathbb{U}}(1)$ that $|\operatorname{{\rm Tr}}T_{z}|=2\Leftrightarrow\tfrac{(\operatorname{{\rm Tr}}T_{z})^{2}}{\det T_{z}}=4$. Moreover, $z\mapsto\tfrac{(\operatorname{{\rm Tr}}T_{z})^{2}}{\det T_{z}}$ is analytic for $z\neq 0$. Therefore, if this would be equal to 4 on an infinite set within the unit circle ${\mathbb{U}}(1)$, then the function would have to be identically $4$, and so would be $\frac{(\operatorname{{\rm Tr}}z^{p}T_{z})^{2}}{\det(z^{p}T_{z})}$. Now, using (1.19), (1.20) one finds for $n\geq 1$
@@
Let us first show Claim 1: For the unique formal solutions (uniqueness follows because it needs to have the correct boundary condition) of ${\mathcal{U}}\Psi=z\Psi$ and ${\mathcal{W}}\Phi=\Psi$ we have that
$\begin{pmatrix}\Phi_{(n,+)}\\ \Psi_{(n,+)}\end{pmatrix}\,=\,T_{z,[0,n]}\begin{pmatrix}1\\ 1\end{pmatrix}\,,\quad\begin{pmatrix}\Phi_{(n,+)}\\ \Psi_{(n,+)}\end{pmatrix}\,=\,T_{z,n}^{\sharp}\,\begin{pmatrix}\Psi_{(n,-)}\\ \Phi_{(n,-})\end{pmatrix}.$
We have the following equivalences:
@@
$\begin{pmatrix}\Phi_{(np,+)}\\ \Psi_{(np,+)}\end{pmatrix}\,=\,T_{z,[0,pn]}\begin{pmatrix}1\\ 1\end{pmatrix}\,=\,T_{z}^{n}T_{z,0}\begin{pmatrix}1\\ 1\end{pmatrix},$
@@
And finally we remark that there is no essential spectrum where $|\operatorname{{\rm Tr}}T_{z}|>2$.
@@
and as $z$ is not an eigenvalue of ${\mathcal{U}}$, Lemma 5.4 implies $y_{z}\neq 0$.Thus, $|y_{z}|>\varepsilon>0$ for some $\varepsilon>0$ and all $z\in e^{i[a,b]}$. Thus, for $f\in C({\mathbb{U}}(1))$ supported in $e^{i[a,b]}$ we have using Theorem 1
@@
Proof of Proposition 1.8.
Note that by Lemma 5.1 the spectrum of ${\mathcal{U}}$ has multiplicity 1 everywhere and is given by the support and measure types of $\mu$. Then, Lemmata 5.3, 5.4 and 5.5 imply that there is no essential spectrum in the interior444interior with respect to the topology on the unit circle ${\mathbb{U}}(1)$. of $\{z\in{\mathbb{U}}(1)\,:\,|\operatorname{{\rm Tr}}T_{z}|\geq 2\}$. So Lemma 5.2 now shows
@@
Together with Lemma 5.4 we deduce that the spectrum is purely absolutely continuous within $\overline{\Sigma}$. Apart from $\overline{\Sigma}$ the spectrum contains only a finite number of eigenvalues outside $\overline{\Sigma}$. Finally, as the Hilbert space is infinite dimensional and the spectrum has multiplicity one, a finite set of eigenvalues can not be the whole spectrum. Thus, $\Sigma\neq\emptyset$. This finishes the proof of Proposition 1.8. ∎
Proof of Theorem 3.
As ${\mathcal{U}}_{\omega}$ is almost surely a compact perturbation of ${\mathcal{U}}$ by condition (C2), we see that as a set, the essential spectrum of ${\mathcal{U}}_{\omega}$ is also given by $\overline{\Sigma}$, and Lemma 5.2 now shows Theorem 3. ∎
@@ Appendix A Facts of linear algebra
$U=\begin{pmatrix}A&B\\ C&D\end{pmatrix}\,\in\,{\mathbb{U}}(n),$
be a unitary $n\times n$ matrix split in blocks of sizes $n_{1},n_{2}$, meaning $A\in{\mathbb{C}}^{n_{1}\times n_{1}}$, $B\in{\mathbb{C}}^{n_{1}\times n_{2}}$, $C\in{\mathbb{C}}^{n_{2}\times n_{1}}$, $D\in{\mathbb{C}}^{n_{2}\times n_{2}}$. Let $P\in{\mathbb{U}}(n_{2})$. Then, the following holds.
@@
    Moreover, if $B$ has trivial kernel then $D-P$ is invertible (for any $P\in{\mathbb{U}}(n_{2})$).
@@
    Assume that $P\in{\mathbb{U}}(n_{2})$ and that $D-P$ is invertible. Then, the Schur complement
    $A-B(D-P)^{-1}C\,\in\,{\mathbb{U}}(n_{1}),$
@@
    $\begin{pmatrix}\mathbf{I}&\mathbf{0}\end{pmatrix}\begin{pmatrix}A&B\\ C&D-P\end{pmatrix}^{-1}\begin{pmatrix}\mathbf{I}\\ \mathbf{0}\end{pmatrix}\,\in\,{\mathbb{U}}(n_{1}),$
@@
    More general, let $P,\,W\in{\mathbb{C}}^{n_{2}\times n_{2}}$ such that $W^{*}P=\mathbf{I}_{n_{2}}$, and $D-P$ and $D-W$ are invertible, then
    $(A-B(D-W)^{-1}C)^{*}(A-B(D-P)^{-1}C)\,=\,\mathbf{I}_{n_{1}}.$
@@
First note that if $(D-P)\varphi=\mathbf{0}$, then
$\|\varphi\|^{2}=\left\|U\begin{pmatrix}\mathbf{0}\\ \varphi\end{pmatrix}\right\|^{2}=\|B\varphi\|^{2}+\|D\varphi\|^{2}=\|B\varphi_{2}\|+\|P\varphi\|^{2}=\|B\varphi\|^{2}+\|\varphi\|^{2},$
@@
In particular, if $B$ has trivial kernel, $\varphi=\mathbf{0}$ and, thus, $D-P$ is invertible.
On the other hand, let $\begin{pmatrix}A&B\\ C&D-P\end{pmatrix}\begin{pmatrix}\varphi_{1}\\ \varphi_{2}\end{pmatrix}=\mathbf{0}$, then
$\|\varphi_{1}\|^{2}+\|\varphi_{2}\|^{2}=\left\|\begin{pmatrix}\varphi_{1}\\ \varphi_{2}\end{pmatrix}\right\|^{2}=\left\|U\begin{pmatrix}\varphi_{1}\\ \varphi_{2}\end{pmatrix}\right\|^{2}=\left\|\begin{pmatrix}\mathbf{0}\\ P\varphi_{2}\end{pmatrix}\right\|^{2}=\|\varphi_{2}\|^{2}.$
Therefore, $\varphi_{1}=\mathbf{0}$ and $D\varphi_{2}=P\varphi_{2}$. Hence,
$\ker\begin{pmatrix}A&B\\ C&D-P\end{pmatrix}\,=\,\left\{\begin{pmatrix}\mathbf{0}\\ \varphi\end{pmatrix}\,:\,\varphi\in\ker(D-P)\;\right\}.$
For part b) note that the first part of b) follows from c). Using the Schur complement identity
$\begin{pmatrix}\mathbf{I}&\mathbf{0}\end{pmatrix}\begin{pmatrix}A&B\\ C&D-P\end{pmatrix}^{-1}\begin{pmatrix}\mathbf{I}\\ \mathbf{0}\end{pmatrix}\,=\,(A-B(D-P)^{-1}C)^{-1},$
@@
and using $W^{*}P=\mathbf{I}$ we obtain
$(A-B(D-W)^{-1}C)^{*}(A-B(D-P)^{-1}C)$
$=A^{*}A-A^{*}B(D-P)^{-1}C-C^{*}(D^{*}-W^{*})^{-1}B^{*}A+C^{*}(D^{*}-W^{*})^{-1}B^{*}B(D-P)^{-1}C$
$=A^{*}A+C^{*}D(D-P)^{-1}C+C^{*}(D^{*}-W^{*})^{-1}D^{*}C+C^{*}(D^{*}-W^{*})^{-1}(\mathbf{I}-D^{*}D)(D-P)^{-1}C$
$=A^{*}A+C^{*}(D^{*}-W^{*})^{-1}\big{(}(D^{*}-W^{*})D+D^{*}(D-P)+\mathbf{I}-D^{*}D\big{)}(D-P)^{-1}C$
$=A^{*}A+C^{*}(D^{*}-W^{*})^{-1}\big{(}(D^{*}-W^{*})(D-P)\big{)}(D-P)^{-1}C=A^{*}A+C^{*}C=\mathbf{I}.$
@@
For $T\in{\mathbb{U}}(1,1)$ we have $|\det T|=1$ and $\tfrac{(\operatorname{{\rm Tr}}T)^{2}}{\det T}\geq 0$. In particular, there is $\chi\in{\mathbb{R}}$ and $\lambda\in{\mathbb{C}}$, such that the eigenvalues of $T$ are given by $e^{i\chi}\lambda$ and $e^{i\chi}\lambda^{-1}$ with $\lambda+\lambda^{-1}\in{\mathbb{R}}$. In particular, $\lambda\in{\mathbb{R}}$ if $|\operatorname{{\rm Tr}}T|=|\lambda+\lambda^{-1}|\geq 2$, and $\lambda\in{\mathbb{U}}(1)$ if $|\operatorname{{\rm Tr}}T|\leq 2$.
@@
then $T^{*}GT=G$ which means $-1=\det(G)=\det(G)|\det T|^{2}$ implying $|\det(T)|=1$. Moreover, we get $|a|^{2}-|c|^{2}=1=|d|^{2}-|b|^{2}$ and $\bar{a}b=\bar{c}d$. On the other hand, $T^{*}GTG=G^{2}=I$ and therefore $TGT^{*}G=I$ giving $TGT^{*}=G$, which gives $|a|^{2}-|b|^{2}=1=|d|^{2}-|c|^{2}$. Hence, $|a|=|d|=\cosh(\gamma)$ and $|b|=|c|=\sinh(\gamma)$ for some $\gamma\geq 0$.
@@
$a=\cosh(\gamma)e^{i\varphi_{1}},\;b=\sinh(\gamma)e^{i\varphi_{2}},\;c=\sinh(\gamma)e^{i\varphi_{3}},\;d=\cosh(\gamma)e^{i\varphi_{4}},$
@@
$e^{-i\varphi_{1}+i\varphi_{2}}\,=\,e^{-i\varphi_{3}+i\varphi_{4}}\quad\text{implying}\quad e^{i\varphi_{1}+i\varphi_{4}}=e^{i\varphi_{2}+i\varphi_{3}}\,.$